Suppose you wanted to use Earth’s magnetic field to make an ac generator at a location where the magnitude of the field is 0.05000 mT. Your coil has 1062 turns and a radius of 3.600 cm. At what angular velocity would you have to rotate it in order to generate an emf of amplitude 1.400 V?
Answer #1

here,

the magnitude of earth magnetic field , B = 0.05 mT = 5 * 10^-5 T

the number of turns , N = 1062 turns

radius , r = 3.6 cm = 0.036 m

let the angular velocity be w

the amplitude of emf , V = N * A * B * w

1.4 = 1062 * pi * 0.036^2 * 5 * 10^-5 * w

solving for w

w = 6.48 * 10^3 rad/s

the angular velocity is 6.48 * 10^3 rad/s